THE ROTATIONAL SPECTRA OF THE WEAKLY BOUND He_{1}-C_{5}H_{5}N, He_{2}-C_{5}H_{5}N, and H_{2}-C_{5}H_{5}N COM-PLEXES

\underline{CHAKREE TANJAROON} and WOLFGANG JÄGER, Department of Chemistry, The University of Alberta, Edmonton, T6G 2G2 Canada.

The rotational spectra of three very weakly bound complexes, He_{1}-C_{5}H_{5}N, He_{2}-C_{5}H_{5}N, and H_{2}-C_{5}H_{5}N, have been studied using a pulsed molecular beam microwave spectroscopic method. Analysis of the rotational and ^{14}N nuclear quadrupole coupling constants of these complexes shows that all three complexes possess the perpendicular geometric configuration. In this configuration, in the dimer, He and H_{2} bind above the aromatic plan of C_{5}H_{5}N and are displaced by less than ten degrees away from the c-axis of the pyridine monomer, toward the nitrogen atom. In the trimer, the second helium atom preferentially binds below the aromatic plan of C_{5}H_{5}N, opposite of the first helium atom. An MP2 ab initio interaction potential between He and C_{5}H_{5}N has been generated and the theoretical results are consistent with those from the experiment. The spectra, molecular structures, quadrupole coupling parameters, and interaction potentials will be discussed.^{a}

aSupport by NSERC and an Alberta Ingenuity Postdoctoral Fellowship to CT are gratefully acknowledged.